A loundspeaker diaphragm is producing a sound for 2.5 s by moving back and forth in a simple harmonic motion. The angular frequency of the motion is 8.22 ✕ 104 rad/s. How many times does the diaphragm move back and forth
Answer #1

Angular frequency = 8.22*10^4 radian /second

Frequency = 8.22*10^4/2*π=1.30*10^4 hertz

Time period =1/ frequency

= 1/1.30*10^4

Total time = number of Period x Time period

2.5. = number of Period ×1/1.30*10^4

Therefore number of Periods

2.5 *1.30*10^4 = number of Periods

32,500 times the diaphragm move back and forth